OpenStudy (anonymous):

Simplify -3x + 9 ----- -3

OpenStudy (anonymous):

(-3x + 9 over -3)

myininaya (myininaya):

\[\frac{-3x+9}{-3}=\frac{-3(x-3)}{-3}\]

myininaya (myininaya):

what is -3/(-3)=?

OpenStudy (anonymous):

divide -3 x by -3 and you get x the divide 9 by -3 and you get -3

myininaya (myininaya):

\[\frac{-3x+9}{-3}=\frac{-3(x-3)}{-3}=\frac{-3}{-3} \cdot (x-3)=1 \cdot (x-3)=x-3\]